# Federated Causal Inference from Multi-Site Observational Data via Propensity Score Aggregation

### v1

@@ 6 Conclusion, limitations and future work
We present a theoretically grounded framework for federated causal inference from decentralized observational data. We introduce two methods for federating local propensity scores—via estimated membership probabilities or density ratios—both of which are designed to be robust to covariate shift and heterogeneity in propensity scores. Our methods improve covariate overlap, leading to more stable and accurate estimates of the ATE. The framework is flexible, allowing the integration of external control arms and supporting heterogeneous local propensity score estimation strategies.
One of the common benefits of federated learning is the ability to increase the overall sample size by leveraging data from multiple sites. While our approach does offer this advantage, it nonetheless requires a sufficiently large amount of data per site, especially when the number of variables is high. Indeed, accurate estimation of key quantities—local propensity scores, parameters of the distribution of the features per site and membership probabilities—relies on sufficient local data. For the latter, the issue becomes more pronounced as the number of sites grows.
RCTs and observational data each have strengths and limitations. Combining them can harness their respective advantages while mitigating weaknesses—supporting goals like improving heterogeneous effect estimation, addressing unobserved confounding, and enhancing generalizability. Estimating treatment effects from diverse sources is thus a promising direction for future federated causal inference. Finally, estimating the CATE in a federated setting is crucial for advancing beyond average treatment effects, towards more personalization in treatment decision-making.
Acknowledgments
This work has been done in the frame of the PEPR SN SMATCH project and has benefited from a governmental grant managed by the Agence Nationale de la Recherche under the France 2030 programme, reference ANR-22-PESN-0003. Rémi Khellaf was supported by an AXIAUM fellowship.
@@ A.2 Proof of Theorem 1
For the AIPW estimator, following the same steps as in section A.2, we have
@@
Because these variances are well defined from what precedes, the Central Limit Theorem can be applied to $\hat{\tau}^{\mathrm{IPW^{*}}}$ and $\hat{\tau}^{\mathrm{AIPW^{*}}}$, which gives the result in Theorem 1. ∎
@@ A.4 Decomposition of the global propensity score
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})e_{k}(X_{i})$ (Eq. 3)
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\frac{P(X_{i}\mid H_{i}=k)}{P(X_{i})}e_{k}(X_{i})$ (Eq. 4)
@@ A.6 Proof of Theorem 4
First, applying Jensen’s inequality with the strictly convex function $t\mapsto\frac{1}{t}$ in ]0;1[]0;1[] 0 ; 1 [ and summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$, we have
@@ A.7 Proof of Theorem 5
Appendix B Federated learning of membership weights with logistic regression
We detail how to estimate membership weights using a multinomial logistic regression model trained via Federated Averaging (FedAvg) McMahan et al., (2017), without requiring access to individual-level data.
Definition 4 (Multinomial Logistic Model).
The multinomial logistic regression model is defined as
$\omega_{k}(X_{i})=\frac{\exp(\theta_{k}^{\top}X_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\theta_{k^{\prime}}^{\top}X_{i})},$ (7)
where $\theta_{k}$ are the parameters of the model for class $k$.
We introduce the following notations: $\Theta=(\theta_{1},\cdots,\theta_{K})\in\mathbb{R}^{d\times K}$ is the matrix containing the parameters of the multinomial logistic regression, and $\ell(\Theta;\mathcal{D})$ is the negative log-likelihood of the model on a dataset $\mathcal{D}$ given by $\ell(\Theta;\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\sum_{k=1}^{K}\mathbb{1}_{\{H_{i}=k\}}\log(\omega_{k}(X_{i}))$ the softmax probabilities. $H_{i}^{\mathrm{enc}}$ is the one-hot encoding of the variable $H_{i}$.
Algorithm 1 presents the FedAvg procedure applied to training a multinomial logistic regression model for estimating membership weights.
Algorithm 1 Federated Multinomial Logistic Regression with FedAvg for Site Membership
1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ rounds of communication, $B$ batch size
2:Initialize model parameters $\Theta_{0}\in\mathbb{R}^{d\times K}$
@@ Appendix B Federated learning of membership weights with logistic regression
5:         $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}^{(k)}$)
@@
7:     $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t}^{(k)}$ // Federated Averaging
@@
11:     $\mathcal{B}_{k}\leftarrow$ a random batch of $B$ samples from $\mathcal{D}_{k}$
12:     $\omega_{l}(x_{i})\leftarrow\frac{\exp(\theta_{l}^{(k)}{}^{\top}x_{i})}{\sum_{l^{\prime}=1}^{K}\exp(\theta_{l^{\prime}}^{(k)}{}^{\top}x_{i})}$ // Multinomial probabilities
13:     $P_{i}\leftarrow(\omega_{1}(X_{i}),\dots,\omega_{K}(X_{i}))$
14:     $\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})\leftarrow\frac{1}{B}\sum_{i\in\mathcal{B}_{k}}x_{i}(P_{i}-H_{i}^{\mathrm{enc}})$ // Multinomial gradient
15:     $\Theta^{(k)}\leftarrow\Theta^{(k)}-\eta\nabla\ell(\Theta^{(k)};\mathcal{B}_{k})$
16:end for
17:return $\Theta^{(k)}$
With a suitable choice of learning rate $\eta$, and a small number of local steps $E$ per round, Algorithm 1 produces parameter estimates for the multinomial logistic regression model that converge to their centralized counterparts as the number of communication rounds $T$ goes to infinity Stich, (2019); Khaled et al., (2020); Li et al., (2019). After $T$ rounds, Algorithm 1 yields an estimate $\widehat{\Theta}^{\mathrm{fed}}=(\hat{\theta}_{1}^{\mathrm{fed}},\cdots,\hat{\theta}_{K}^{\mathrm{fed}})$ of the multinomial logistic parameters which can be used to estimate the membership weights as:
$\hat{\omega}_{k}(X_{i})=\frac{\exp(\hat{\theta}_{k}^{\mathrm{fed}}{}^{\top}x_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\hat{\theta}_{k^{\prime}}^{\mathrm{fed}}{}^{\top}x_{i})}.$
@@ Appendix C Simulation Details
[…]
@@
$\gamma_{k}$ [.$\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$ No overlap: not logistic (only controls) Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$ Good overlap: $\gamma_{2}^{\mathrm{(good)}}$ [.$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
@@
$\mu_{k}$ $(1,\dots,1)\in\mathbb{R}^{d}$ $(1.5,\dots,1.5)\in\mathbb{R}^{d}$ $(3,\dots,3)\in\mathbb{R}^{d}$
@@ Appendix D Overlap Improvement
Echoing theorem 5, Figures 5 and 6 display the empirical distributions (on a log-scale) of the local propensity score in site 2 ($e_{2}$) and of the global propensity score $e$ in the Poor local overlap scenario (see Section 5 for corresponding results). A good overlap is when both the propensity score distributions for the treated and control overlap, and are far from 0. We see that the poor overlap at site 2 (Figure 5), with values of $e_{2}$ on the local data close to 0, is significantly improved at the global level (Figure 6).
@@
Figure 6: Global overlap for the Poor local overlap scenario (DGP A). We see a clear improvement compared to the local overlap in site 2 (Figure 5).

### v2

@@ 5 Conclusion, limitations and future work
We propose a theoretically grounded framework for federated causal inference that leverages membership weights to construct valid pseudo-populations across silos. These weights, estimated via flexible parametric or nonparametric models, improve overlap and yield more stable ATE estimates, without sharing raw data. Our framework accommodates heterogeneous local propensity score estimation strategies, supports external control arms, and remains robust to even extreme local treatment–control imbalances. Although sufficiently large per-site datasets are still needed—particularly in high-dimensional settings—our approach is especially well suited to a moderate number of large silos, where FL most effectively increases effective sample size.
Promising avenues for future work include principled handling of site effects (i.e., relaxing Assumption 2) and extending our framework to CATE estimation. While the ATE remains central in econometrics, biomedicine, and public policy, considering CATE is important for moving towards personalization. Our work lays a foundation for federated CATE estimation: most learners (T-, S-, X-, and R-learners) rely on nuisance quantities such as propensity scores, and our federated estimation framework could be incorporated into these methods, although further work is needed to assess its formal properties and practical performance.
@@ A.2 Proof of Theorem 1
For the AIPW estimator, following the same steps as in section˜A.2, we have
@@
Because these variances are well defined from what precedes, the Central Limit Theorem can be applied to $\hat{\tau}^{\mathrm{IPW^{*}}}$ and $\hat{\tau}^{\mathrm{AIPW^{*}}}$, which gives the result in Theorem˜1. ∎
@@ A.4 Decomposition of the global propensity score
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k)\frac{P(X_{i}\mid H_{i}=k)}{P(X_{i})}e_{k}(X_{i})$ (DW)
$=\sum_{k=1}^{K}\mathbb{P}(H_{i}=k\mid X_{i})e_{k}(X_{i})$ (MW)
@@ A.6 Proof of Theorem 4
First, applying Jensen’s inequality with the strictly convex function $t\mapsto\frac{1}{t}$ in ]0;1[]0;1[ and summing-to-one weights $\omega_{k}(X)=\mathbb{P}(H_{i}=k\mid X_{i}=X)$, we have
@@ A.7 Proof of Theorem 5
Appendix B Federated learning of membership weights
We describe how to estimate membership weights using a general parametric classification model trained via Federated Averaging (FedAvg) [McMahan et al., 2017], without requiring access to individual-level data.
B.1 General Parametric Model
Let $\omega_{k}(X_{i};\Theta)$ denote the predicted membership probability of site $k$ for covariate vector $X_{i}$, where $\Theta$ are the parameters of the model. For a generic parametric model, $\Theta\in\mathbb{R}^{p}$ represents all trainable parameters, and the membership-weight estimation problem can be formulated as the minimization of the empirical cross-entropy loss:
$\ell(\Theta;\mathcal{D})=-\frac{1}{|\mathcal{D}|}\sum_{i=1}^{|\mathcal{D}|}\sum_{k=1}^{K}\mathbb{1}_{\{H_{i}=k\}}\log\bigl(\omega_{k}(X_{i};\Theta)\bigr),$
where $H_{i}^{\mathrm{enc}}$ is the one-hot encoding of the site membership $H_{i}$.
This framework encompasses a broad class of models, including:
  • Multinomial logistic regression:
    $\omega_{k}(X_{i};\Theta)=\frac{\exp(\theta_{k}^{\top}X_{i})}{\sum_{k^{\prime}=1}^{K}\exp(\theta_{k^{\prime}}^{\top}X_{i})},\quad\Theta=(\theta_{1},\dots,\theta_{K})\in\mathbb{R}^{d\times K}.$
  • Neural networks, where $\Theta$ includes all weights and biases across layers, and $\omega_{k}(\cdot)$ is the softmax output of the final layer.
B.2 Federated Training Procedure
Algorithm 1 summarizes the generic Federated Averaging procedure for training any differentiable parametric model for membership-weight estimation.
Algorithm 1 Federated Learning of Membership Weights with a Parametric Model
1:Input: $K$ sites, $E$ local steps, $\eta$ learning rate, $T$ communication rounds, $B$ batch size
2:Initialize global parameters $\Theta_{0}$
@@ B.2 Federated Training Procedure
5:   $\Theta_{t+1}^{(k)}\leftarrow$ LocalUpdate($k,\Theta_{t}$)
@@
7:  $\Theta_{t+1}\leftarrow\sum_{k=1}^{K}\frac{n_{k}}{n}\Theta_{t+1}^{(k)}$ // FedAvg aggregation
@@
11:  Sample mini-batch $\mathcal{B}_{k}$ of size $B$ from local data $\mathcal{D}_{k}$
12:  Compute predictions $\omega_{k}(X_{i};\Theta)$ for $i\in\mathcal{B}_{k}$
13:  Compute gradient $\nabla\ell(\Theta;\mathcal{B}_{k})$ of cross-entropy loss
14:  $\Theta\leftarrow\Theta-\eta\nabla\ell(\Theta;\mathcal{B}_{k})$
15:end for
16:return $\Theta$
With a suitable choice of learning rate $\eta$ and a moderate number of local steps $E$ per round, Algorithm 1 converges to the same solution as centralized training as the number of communication rounds $T\to\infty$ [Stich, 2019, Khaled et al., 2020, Li et al., 2019]. The resulting federated parameter estimate $\widehat{\Theta}^{\mathrm{fed}}$ yields the estimated membership weights:
$\hat{\omega}_{k}(X_{i})=\omega_{k}(X_{i};\widehat{\Theta}^{\mathrm{fed}}),\qquad k=1,\dots,K.$
B.3 Example: Multinomial Logistic Regression
When the model is multinomial logistic regression, $\Theta$ is the matrix of regression coefficients, and the local gradient takes the closed form:
$\nabla\ell(\Theta;\mathcal{B}_{k})=\frac{1}{B}\sum_{i\in\mathcal{B}_{k}}X_{i}\bigl(\omega(X_{i};\Theta)-H_{i}^{\mathrm{enc}}\bigr),$
where $\omega(X_{i};\Theta)$ is the vector of softmax probabilities for sample $i$.
@@ Appendix C Simulation Details
The parameters common to all settings are shown in Table 1, where $\gamma_{2}^{\mathrm{(weak)}}=[-2.5,-1,-0.15,-0.15,0,-0.15,-1,-0.15,-0.15,0]$ and $\gamma_{2}^{\mathrm{(good)}}=[-.05,-.1,-.05,-.1,.05,-.1,-.05,-.1,.05,-.1]$.
@@
$\gamma_{k}$ $\begin{array}[]{c}[-.25,.25,\\ -.25,-.25,.25,\\ -.25,-.25,.25,\\ -.25,.25]\end{array}$
No overlap: not logistic
(only controls)
Weak overlap: $\gamma_{2}^{\mathrm{(weak)}}$
Good overlap: $\gamma_{2}^{\mathrm{(good)}}$
$\begin{array}[]{c}[.15,-.15,\\ .15,-.15,.15,\\ -.15,.15,-.15,\\ .15,-.15]\end{array}$
@@
$\mu_{k}$ $(1,\dots,1)\in\mathbb{R}^{d}$ $(1.5,1.5,1.5,1,\dots,1)\in\mathbb{R}^{d}$ $(2,2,2,1,\dots,1)\in\mathbb{R}^{d}$
@@ Appendix D Overlap Improvement
Echoing theorem˜5, Figures 4 and 5 display the empirical distributions (on a log-scale) of the local propensity score in site 2 ($e_{2}$) and of the global propensity score $e$ in the Poor local overlap scenario (see Section˜4 for corresponding results). A good overlap is when both the propensity score distributions for the treated and control overlap, and are far from 0. We see that the poor overlap at site 2 (Figure 4), with values of $e_{2}$ on the local data close to 0, is significantly improved at the global level (Figure 5).
@@
Figure 5: Global overlap for the Poor local overlap scenario (DGP A). We see a clear improvement compared to the local overlap in site 2 (Figure˜4).
Appendix E Additional Simulations
E.1 Large number of sites
We further investigate the robustness of our estimators as the number of sites increases. Specifically, we simulate $K=20$ sites under weak local overlap (corresponding to the setting of Figure 4 in the main paper) and report the bias and mean squared error (MSE) for both data-generating processes (DGP A and DGP B).
Table 4: Bias and mean squared error (MSE) under $K=20$ sites with weak local overlap. Bold values indicate unbiasedness or lowest MSE.
Method Bias (DGP A) MSE (DGP A) Bias (DGP B) MSE (DGP B)
Oracle IPW 0.00 0.298 0.00 0.130
Fed. IPW-MW (logistic) 0.48 0.294 0.00 0.095
Fed. IPW-MW (NN) 0.00 0.240 0.00 0.123
Fed. IPW-DW 0.00 0.096 0.50 0.450
Meta-SW IPW -2.09 13.279 -1.77 7.977
Oracle AIPW 0.00 0.007 0.00 0.049
Fed. AIPW-MW (logistic) 0.79 0.639 0.00 0.078
Fed. AIPW-MW (NN) 0.00 0.153 0.00 0.084
Fed. AIPW-DW 0.00 0.019 0.25 0.194
Meta-SW AIPW -1.24 1.775 -2.36 18.465
As shown in Table 4, estimators using neural network membership weights consistently achieve unbiasedness across both DGPs, although with slightly higher variance compared to their well-specified parametric counterparts. This is consistent with the behavior of more flexible models, which trade off a small increase in variance for robustness to misspecification. When parametric assumptions are correctly specified (e.g., Gaussian density weights in DGP A, logistic membership weights in DGP B), these methods achieve the lowest MSE. Overall, the NN-based approach demonstrates robust consistency across DGPs and scales well with the number of sites.
E.2 Fully model-agnostic ATE estimation
We additionally investigate the benefits of fully model-agnostic estimation for membership weights $P(H\mid X)$. In this setting, we use federated neural networks to estimate membership probabilities in a flexible way and pair them with non-parametric local propensity score estimators (e.g., Random Forests). This procedure does not require prior parametric knowledge about $H\mid X$ or $X\mid H$, making it robust to a wide range of DGPs.
We simulate $K=3$ sites with $d=10$ covariates, $n_{k}=600$ observations per site, and $X\mid H=k\sim D_{k}$ as a mixture of Gaussians with weak local overlap. The results are summarized in Table 5.
Method Bias MSE
Oracle IPW 0.00 0.116
Fed. IPW-MW (NN MW + RF $e_{k}$) 0.00 0.075
Meta-SW IPW (RF $e_{k}$) 0.89 0.884
Oracle AIPW 0.00 0.030
Fed. AIPW-MW (NN MW + RF $e_{k}$) 0.00 0.040
Meta-SW AIPW (RF $e_{k}$) -0.36 0.180
Table 5: Bias and MSE under fully non-parametric estimation of membership weights and local propensity scores ($K=3$, weak local overlap). Bold values indicate unbiased (or nearly unbiased) estimators.
As shown in Table 5, our federated approach with neural-network-based membership weights and random-forest local propensities yields nearly unbiased estimates with substantially lower MSE compared to meta-analysis estimators. This demonstrates that misspecifications in either the membership weights or the local propensity scores can be effectively mitigated by adopting fully non-parametric methods in a federated learning setting.
E.3 Robustness to local misspecifications
Similar to meta-analysis estimators, we argue that inconsistencies in local propensity models can be mitigated as $K$ grows, provided that such inconsistencies are marginal relative to the number of correctly specified sites. To illustrate this, we simulate $K$ sites under DGP B, each holding $n_{k}=60$ observations drawn from homogeneous covariate distributions and sharing the same true propensity model (logistic). We then intentionally misspecify the propensity model for site $k=1$, setting $\widehat{e}_{1}(x)\equiv 0.1$ for all $x$.
Table 6 reports the bias and mean squared error (MSE) of several estimators under two scenarios: $K=3$ (few sites) and $K=200$ (many sites).
Method Bias ($K=3$) MSE ($K=3$) Bias ($K=200$) MSE ($K=200$)
Oracle IPW 0.00 0.272 0.00 0.002
Fed. IPW (logistic MW) 0.38 0.344 0.00 0.002
Meta IPW 2.61 7.902 0.03 0.003
Oracle AIPW 0.00 0.058 0.00 0.001
Fed. AIPW (logistic MW) -0.33 0.325 -0.01 0.002
Meta AIPW -0.51 0.481 -0.07 0.007
Table 6: Bias and mean squared error (MSE) under local misspecification of site $k=1$ propensity model.
With only $K=3$ sites, both federated and meta-analysis estimators exhibit substantial bias. However, as the number of sites increases, the influence of the single misspecified site becomes negligible, and the bias decreases markedly (e.g., from $0.38$ to $0.00$ for Federated IPW with logistic membership weights). As expected, the AIPW estimator retains its double-robustness property even under misspecification of the outcome model, leading to consistently low bias and MSE across settings.